# Geometric theory of quark confinement

### v5

@@ 8 Conclusions
We have derived the high levels of $\bar{q}q$ meson mass spectrum associated with the confining area law in the quenched approximation. By solving the spectral equation for the Hodge-dual minimal surface, we obtain a linear Regge trajectory $M^{2}\sim J$, consistent with the area law of confinement. Our confinement mechanism does not provide masses to glueballs; this would require further investigation.
@@
The spacetime dimensionality $D=4$ plays the unique role in this solution. The confinement mechanism relies fundamentally on the self-duality of the area derivative—a geometric property that exists exclusively in four dimensions. This stands in sharp contrast to fundamental bosonic string models, which require unphysical critical dimensions (e.g., $D=26$) to maintain quantum consistency. While effective string theories can describe long flux tubes in $D=4$, our result shows that the fundamental confining mechanism is a stable, non-vibrating minimal surface, exempt from the quantization anomalies of the fundamental Nambu-Goto string.
@@
Our solution demonstrates that the confining surface is stable specifically in $D=4$. This dimensionality constraint explains the absence of daughter trajectories in the resulting spectrum: the confining object is a minimal surface fixed by 4D geometry, not a fundamental string vibrating in higher embedding dimensions. The “missing” string modes are eliminated, leaving a clean linear Regge trajectory consistent with phenomenological observation.
Acknowledgements.—The author thanks colleagues at the Institute for Advanced Study for inspiring discussions.
Appendix A General Theory: Spinor Factorization and Analytical Structure
The mathematical problem of minimization of the Dirichlet functional with Virasoro constraint in four dimensions is a well known problem in modern mathematics. Let us briefly summarize the existing theory.
A.1 Twistor Framework and the Klein Quadric
The fundamental constraint of the theory is the null condition on the holomorphic tangent vector of the minimal surface, $(f^{\prime})^{2}=\sum_{\mu=1}^{4}(f^{\prime}_{\mu})^{2}=0$. Geometrically, this defines the Klein Quadric in the complexified tangent space $\mathbb{CP}^{3}$. Using the isomorphism $SO(4,\mathbb{C})\cong SL(2,\mathbb{C})_{L}\times SL(2,\mathbb{C})_{R}$, we map the vector $f^{\prime}_{\mu}$ to a $2\times 2$ matrix $f^{\prime}_{a\dot{b}}=f^{\prime}_{\mu}(\sigma_{\mu})_{a\dot{b}}$. The null condition corresponds to $\det(f^{\prime})=0$, ensuring the matrix has rank 1. Thus, the tangent vector factorizes globally into a product of two spinors:
$f^{\prime}_{a\dot{b}}(z)=\lambda_{a}(z)\mu_{\dot{b}}(z)$ (A.1)
where $\lambda_{a}$ and $\mu_{\dot{b}}$ are meromorphic sections of spinor bundles over the worldsheet. This factorization linearizes the quadratic Virasoro constraint, replacing it with the problem of determining the analytic properties of the spinors.
A.2 Gauge Invariance and Birkhoff Factorization
The physical vector $f^{\prime}$ is invariant under the local complex gauge transformation:
$\lambda_{a}(z)\to w(z)\lambda_{a}(z),\qquad\mu_{\dot{b}}(z)\to w^{-1}(z)\mu_{\dot{b}}(z)$ (A.2)
where $w(z)$ is a meromorphic function. This redundancy is governed by the Birkhoff-Grothendieck theorem on the splitting of vector bundles over $\mathbb{P}^{1}$. The gauge function $w(z)$ allows us to distribute the zeroes and poles between $\lambda$ and $\mu$, classifying solutions by integer topological indices (partial indices) related to the winding number of the gauge field required to regularize the spinors at infinity.
A.3 Analytic Structure via Plemelj Projection
The connection to the physical loop $C(\sigma)$ (where $\sigma=e^{\mathrm{i}\theta}$) is achieved via the Plemelj-Sokhotski decomposition. We expand the loop coordinates into positive and negative frequency modes:
$C(\sigma)=C_{+}(\sigma)+C_{-}(\sigma)=\sum_{n\geq 0}c_{n}\sigma^{n}+\sum_{n<0}c_{n}\sigma^{n}$ (A.3)
If the loop is algebraic, $C_{-}(\sigma)$ extends to a rational function $C_{-}(z)$ in the complex plane $|z|>1$. The poles of the spinor solution $f^{\prime}(z)$ are dictated by the singularities of $C_{-}(z)$. Analytically, the problem is to find the "closest" null vector $f^{\prime}$ to the derivative $\partial_{z}C_{+}$. This is equivalent to finding a subspace $W$ in the Sato Grassmannian that satisfies the isotropic condition and matches the boundary data.
A.4 Explicit Algebraic Solutions
The general theory of the spinor factorization predicts that rational choices for the sections $\lambda_{a}(z)$ and $\mu_{\dot{b}}(z)$ generate algebraic minimal surfaces. These correspond to "Finite Gap" solutions where the algebraic curve has a finite genus. The distribution of poles between the spinors determines the topology and boundary behavior of the surface.
We illustrate this with four classical examples, reinterpreted in our spinor language: 1. The Enneper Surface (Polynomial Solution). This is the simplest solution with trivial topology (genus zero, one boundary at infinity), corresponding to polynomial spinors with a single pole at $z=\infty$.
$\lambda(z)=\begin{pmatrix}1\\ z\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}1\\ -z\end{pmatrix}$ (A.4)
The resulting null vector $f^{\prime}(z)\sim(1-z^{2},i(1+z^{2}),2z,0)$ integrates to a cubic polynomial map. This surface, discovered by Enneper in 1864 [enneper1864], represents the fundamental "ground state" of the algebraic solutions. 2. The Catenoid (Rational Solution with Poles). To generate a surface with the topology of an annulus (the only minimal surface of revolution), one must introduce simple poles into the spinors.
$\lambda(z)=\begin{pmatrix}z^{-1/2}\\ z^{1/2}\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}z^{-1/2}\\ -z^{1/2}\end{pmatrix}$ (A.5)
Using the gauge freedom $\lambda\to w\lambda,\mu\to w^{-1}\mu$ with $w=z^{-1/2}$, we can redistribute the poles to find the standard rational form. Euler [euler1744] originally proved the minimality of this surface, which corresponds to the physical shape of a soap film stretched between two rings. 3. The Henneberg Surface (Non-Orientable). A more complex rational ansatz generates the Henneberg surface [henneberg1875], which is a non-orientable minimal surface (a realization of the projective plane in 3D space). It arises from a rational map of degree 4:
$\lambda(z)=\begin{pmatrix}1-z^{-2}\\ z-z^{-1}\end{pmatrix},\qquad\mu(z)=\begin{pmatrix}1\\ -1\end{pmatrix}$ (A.6)
4. The Helicoid (Transcendental Solution). The Helicoid is the locally isometric conjugate surface to the Catenoid. In the spinor formalism, it arises from the same meromorphic data but with a phase shift that exposes the period of the complex logarithm.
$\lambda(z)=e^{i\pi/4}\begin{pmatrix}z^{-1/2}\\ z^{1/2}\end{pmatrix},\qquad\mu(z)=e^{i\pi/4}\begin{pmatrix}z^{-1/2}\\ -z^{1/2}\end{pmatrix}$ (A.7)
The resulting null vector possesses a pole with purely imaginary residue, $f^{\prime}(z)\sim i/z$. Integration yields a logarithmic branch cut $f(z)\sim i\ln z$, corresponding to the multi-valued height function of the double helix ($X_{3}\propto\Im(\ln z)=\theta$).

### v6

@@ 8 Conclusions
We have established the geometric foundation for the string dual of Planar QCD. The confinement mechanism relies fundamentally on the self-duality of the area derivative—a geometric property that exists exclusively in four dimensions. This stands in sharp contrast to fundamental bosonic string models, which require unphysical critical dimensions (e.g., $D=26$) to maintain quantum consistency. While effective string theories can describe long flux tubes in $D=4$, our result shows that the fundamental confining mechanism is a stable, non-vibrating minimal surface, exempt from the quantization anomalies of the fundamental Nambu-Goto string.
@@
Our results demonstrate that the confining surface is stable specifically in $D=4$. This dimensionality constraint explains the absence of daughter trajectories in the resulting spectrum: the confining object is a minimal surface fixed by 4D geometry, not a fundamental string vibrating in higher embedding dimensions. The “missing” string modes are eliminated, leaving a clean linear Regge trajectory consistent with phenomenological observations.
In the companion paper [16], we populate this rigid surface with internal Majorana fermions. We show that this "Fermi String" theory reproduces the planar graphs of QCD and leads to a finite effective action in Momentum Loop Space, allowing for the computation of the meson spectrum.
Acknowledgements
The author is grateful to Nima Arkani-Hamed for the invitation to present this work at the IAS Particle Theory "Pizza" Seminar, which provided a stimulating environment for discussing these results. I also thank Edward Witten for valuable discussions regarding the distinction between this confining geometry and the conventional area law.